# Multi-objective Aerial IRS-assisted ISAC Optimization via Generative AI-enhanced Deep Reinforcement Learning

### v1

@@ III-B2 Reflected Link
where $\alpha^{\text{br}}$ in the BS-aerial IRS link, and $d^{\text{br}}[t]=\sqrt{\|q^{\text{b}}-q^{\text{r}}[t]\|^{2}}$ represents the distance from the BS to the aerial IRS at time slot $t$. Moreover, the LoS component $g_{\text{L}}^{\text{br}}[t]$ can be denoted as
@@
where $d_{r}$ and $d_{s}$ represent the spacing between IRS elements and the separation between neighboring antennas at the BS, respectively. Moreover, $\lambda$ is the wavelength. In addition, $\zeta$ and $\iota$ denote the angles of arrival (AoA) in the vertical and horizontal directions at the aerial IRS, respectively.
Likewise, the channel vector $h_{n}^{\text{ru}}[t]\in\mathbb{C}^{L\times 1}$ represents the communication link between the aerial IRS and the $n$-th user in time slot $t$, which is given by
@@
where $\alpha^{\text{ru}}$ indicates the path loss. Moreover, $d_{n}^{\text{ru}}[t]=\sqrt{\|q^{\text{r}}[t]-q_{n}^{\text{u}}\|^{2}}$ is used to represent the distance from the aerial IRS and the $n$-th user. In addition, the LoS component $g_{n,\text{L}}^{\text{ru}}[t]$ is given by
@@
where $\zeta_{n}$ is the angle of departure (AoD) from the aerial IRS to the $n$-th user.
III-B3 Composite Link
@@
where $\mathbf{v}[t]=\left[e^{j\theta_{1}[t]},e^{j\theta_{2}[t]},\cdots,e^{j\theta_{L}[t]}\right]$ is the passive beamforming vector of the IRS, and $\theta_{l}[t]\in[0,2\pi)$ is the phase shift of the $l$-th IRS element, wherein $l=\{1,2,\cdots,L\}$.
@@
III-C Communication Model
Let $s_{n}[t]$ and $\omega_{n}[t]\in\mathbb{C}^{M\times 1}$ denote the symbol transmitted by the BS to the $n$-th user and the corresponding active beamforming vector in time slot $t$, respectively. We assume the symbol are independent and identically distributed with zero mean and unit variance, expressed as $s_{n}\sim\mathcal{N}(0,1)$. Therefore, the BS transmission signal can be represented by
@@ III-C Communication Model
where $\varrho_{n}[t]\sim\mathcal{N}(0,\sigma_{n}^{2})$ represents the additive white Gaussian noise characterized by a variance $\sigma_{n}^{2}$ in time slot $t$.
@@
$\gamma_{n}[t]=\frac{\left|\widetilde{H}_{n}^{\text{bu}}[t]\omega_{n}[t]\right|^{2}}{\sum_{i=1,i\neq n}^{N}\left|\widetilde{H}_{n}^{\text{bu}}[t]\omega_{i}[t]\right|^{2}+\sigma_{n}^{2}}.$ (11)
III-D Sensing Model
Notably, ISAC can use the transmitted communication signals for sensing operations by sharing the spectrum resources [12], thereby achieving efficient resource utilization. Given that the presence of obstacles impacts the sensing performance of the BS, an aerial IRS is deployed to reconfigure the sensing link by adjusting the phase of its reflective elements, enabling the direct target sensing services [22]. Thus, the steering vector from the IRS to the target can be expressed as [10]
$a[t]=\alpha[t]\left[1,e^{j\frac{2\pi}{\lambda}\sin({\vartheta^{\text{ST}}}[t])},\cdots,e^{j\frac{2\pi(L-1)}{\lambda}\sin({\vartheta^{\text{ST}}}[t])}\right]^{T},$ (12)
where $\vartheta^{\text{ST}}[t]$ is the angle of interest in time slot $t$, and $\alpha[t]$ represents the large-scale fading coefficient.
Then, the received signal of the target can be denoted as [10]
$\widetilde{\mathbf{x}}[t]=a^{\text{H}}[t]\Phi[t]h^{\text{br}}[t]\left(\sum_{n=1}^{N}\omega_{n}[t]s_{n}[t]\right)+\varrho^{\text{ST}}[t],$ (13)
where $\varrho^{\text{ST}}[t]\sim\mathcal{N}(0,\sigma_{\text{g}}^{2})$ is the noise received at the target. The beampattern gain of the IRS for the target in time slot $t$ is defined as
@@ III-D Sensing Model
$=\mathbf{g}^{\text{H}}[t]\left(\sum_{n=1}^{N}\omega_{n}[t]\omega_{n}^{\text{H}}[t]\right)\mathbf{g}[t],$
where $\mathbf{g}^{\text{H}}[t]=a^{\text{H}}[t]\Phi[t]h^{br}[t]$.
In this case, the echo rate of the target can be represented as [12]
$R^{\text{ST}}[t]=\log_{2}(1+\frac{\mathbf{G}[t]}{\sigma_{\text{g}}^{2}}).$ (15)
III-E UAV Mobility and Energy Model
We assume that the UAV flies at a fixed altitude $z^{\text{r}}$ within the target area for task execution. The horizontal location $\left(x^{\text{r}}[t],y^{\text{r}}[t]\right)$ of the UAV (i.e., aerial IRS) in time slot $t$ depends on the flight speed $v_{\text{u}}[t]$ and yaw angle $\theta_{\text{u}}[t]\in[-\pi,\pi]$, which can be denoted as
@@ III-E UAV Mobility and Energy Model
Due to the limited energy resources of the UAV, optimizing propulsion energy consumption can enhance endurance, thereby extending mission execution time and improving the overall operational efficiency of the system. The propulsion energy of the UAV in time slot $t$ can be expressed as
@@
IV Problem Formulation and Analyses
In this section, we aim to formulate CSEMOP that focuses on the communication, sensing, and energy efficiency performance of the considered aerial IRS-assisted ISAC system. We first introduce the optimization variables and objectives. Based on this, the optimization problem is formulated and analyzed.
IV-A Problem Formulation
@@
Objective 2: Echo rate is a common metric for evaluating the sensing performance in an ISAC system. Thus, the second objective aims to maximize the echo rate of the target, which is given by
@@
Objective 3: UAV is an energy-constrained airborne platform, and reducing its energy consumption can extend operating time of the considered system. Therefore, the third objective focuses on minimizing propulsion energy, expressed as follows:
@@
$-\pi\leq\theta_{\rm u}[t]\leq\pi,\quad\forall t\in\mathcal{T},$ (21e)
@@
where the constraints (21b) and (21c) represent the boundaries of the area within which the UAV performs the mission. Moreover, the flight speed and yaw angle in each time slot are bounded by constraints (21d) and (21e). In addition, constraint (21f) indicates that the transmission power of the BS is less than $P_{\rm max}$. Furthermore, constraint (21g) specifies the phase shift constraint of the IRS.
IV-B Problem Properties
We can find that the formulated CSEMOP has the following natures:
Non-convexity: According to the three optimization objectives presented in Eqs. (18-20), it can be observed that the optimization variables $\mathbf{\Omega}$, $\mathbf{\Psi}$, $\mathbf{\mathcal{V}}$, and $\mathbf{\Theta}$ are coupled. Moreover, IRS phase shift $\theta_{l}$[t] appears exponentially in the IRS beamforming vector [16]. As a result, CSEMOP is highly non-convex, making it difficult to solve.
@@ IV-B Problem Properties
Trade-off: It can be observed that there are conflicts among the three optimization objectives. For example, focusing on the communication performance between the BS and the users will impact the sensing capability of the BS for the target under given channel conditions. Moreover, improving both sensing and communication performance of the system requires rapid deployment of the aerial IRS to suitable regions, which increases UAV propulsion energy consumption. Therefore, there is a trade-off among the three optimization objectives.
V Proposed Algorithm
In this section, we first clarify the motivation for using DRL to solve CSEMOP. Subsequently, we reformulate the CSEMOP as a MDP. Next, we present the core principles of the conventional DDPG algorithm and the improved factors. Finally, we show the proposed GDMDDPG method and analyze the computation complexity.
V-A Motivation of Using DRL
For traditional optimization algorithms, such as convex optimization, the optimization problem are typically decomposed into several subproblems and solved using alternating optimization algorithms, which may influence the accuracy of the final result. Moreover, aerial IRS makes the system highly dynamic, and the prior knowledge required by traditional optimization algorithms is difficult to obtain in such a dynamic environment. In addition, the considered system evolves over time, making it a significant challenge to balance short-term and long-term gains of the system during task execution. However, traditional optimization algorithms have limitations in achieving such balance.
In contrast, DRL algorithms use the trial-and-error learning mechanism through interactions between the agent and the environment, and possess superior capabilities for balancing short-term and long-term rewards in the dynamic environment, providing a feasible solution to the aforementioned challenges. Therefore, we propose a GDMDDPG method to solve the considered CSEMOP.
V-B MDP Construction
@@
V-B1 State Space $\mathcal{S}$
The environment state $s[t]\in\mathcal{S}$ observed by the agent in time slot $t$ is defined as $s[t]=\{t,q^{r}[t]\}$, where $t$ represents the index of current time slot in the pre-defined time slot set $\mathcal{T}$, and $q^{r}[t]$ is the location of the aerial IRS in time slot $t$.
V-B2 Action Space $\mathcal{A}$
According to the observed environment state $s[t]$ in time slot $t$, the agent cantake the action $a[t]\in\mathcal{A}$, which is denoted as $a[t]=\{\omega[t],\phi[t],v_{u}[t],\theta_{u}[t]\}$. Specifically, $\omega[t]=\{\omega_{n}[t]|n\in\mathcal{N}\}$ represents the beamforming matrix of the BS for all users, and $\phi[t]$ represents the phase shifts of the aerial IRS in time slot $t$. Moreover, $v_{u}[t]$ and $\theta_{u}[t]$ denote the flight velocity and yaw angle of the UAV in time slot $t$.
V-B3 Reward Function $\mathcal{R}$
@@
V-C Principles of DDPG Algorithm
@@
V-C1 Actor-Critic Framework
@@
V-C2 Experience Replay Buffer
@@ V-C3 Network Update
To avoid significant changes in the target network during training, which could lead to instability in the agent learning process, the target networks are updated using a soft-update method as follows:
@@ V-D GDMDDPG
V-D1 The Limitations of Conventional DDPG
@@
First, in conventional DDPG algorithm, the actor network typically employs MLP as the core architecture due to the simplicity and ease of implementation. However, in more complex environments, the fixed MLP architecture may fails to fully capture and analyze the intricate information present in the environment states, which in turn limits the decision-making capabilities of the agent. Notably, diffusion model provides a viable solution to address the above limitations due to its strong ability in modeling complex data distributions and generating high-quality samples [35]. By replacing the traditional MLP with the diffusion model, the actor network can more effectively analyze the environment states, thereby generating superior actions and further enhancing the overall system performance.
@@
Second, DDPG is based on deterministic policy, where the actor network directly outputs a specific action value rather than a probability distribution over actions. Such a deterministic policy limits the agent ability to explore the action space effectively. To address this issue, noise can be added to the actions generated by the actor network, introducing randomness to promote the agent exploration of the action space. Moreover, this noise perturbation mechanism helps alleviate the tendency of the DDPG to output action boundary values during early training phase, thereby enhancing the convergence speed of the training process.
@@
Third, the experience replay buffer in conventional DDPG employs a random sampling strategy, which might fail to prioritize experiences that hold significant learning value, resulting in slower convergence. To address this issue, the RPER sampling mechanism, which integrates ERE and PER techniques, can be incorporated into the network update process of the DDPG. Specifically, ERE technique ensures that more recent experiences are sampled more actively. Based on this, PER technique assigns a priority to each experience sample, allowing those that contribute more to learning to be sampled with higher probability. As such, combining PER and ERE can effectively mitigate the limitations of the random sampling strategy.
@@
In the following, we present an in-depth explanation of the diffusion model-based actor network, noise perturbation mechanism, and RPER sampling mechanism.
V-D2 Diffusion Model-based Actor Network

### v2

@@ 3.2.2 Reflected Link
where $\alpha^{\text{br}}$ is the path loss in the BS-IRS link, and $d^{\text{br}}[t]=\sqrt{\|q^{\text{b}}-q^{\text{r}}[t]\|^{2}}$ represents the distance from the BS to the low-altitude IRS in time slot $t$. Moreover, the LoS component $g_{\text{L}}^{\text{br}}[t]$ can be denoted as
@@
where $d_{r}$ and $d_{s}$ represent the spacing between IRS elements and the separation between neighboring antennas at the BS, respectively. Moreover, $\lambda$ is the wavelength. In addition, $\zeta$ and $\iota$ denote the angles of arrival (AoA) in the vertical and horizontal directions at the low-altitude IRS, respectively.
Likewise, the channel vector $h_{n}^{\text{ru}}[t]\in\mathbb{C}^{L\times 1}$ represents the communication link between the low-altitude IRS and the $n$-th user in time slot $t$, which is given by
@@
where $\alpha^{\text{ru}}$ indicates the path loss. Moreover, $d_{n}^{\text{ru}}[t]=\sqrt{\|q^{\text{r}}[t]-q_{n}^{\text{u}}\|^{2}}$ is used to represent the distance from the low-altitude IRS and the $n$-th user. In addition, the LoS component $g_{n,\text{L}}^{\text{ru}}[t]$ is given by
@@
where $\zeta_{n}$ is the angle of departure (AoD) from the low-altitude IRS to the $n$-th user.
3.2.3 Composite Link
@@
where $\mathbf{v}[t]=\left[e^{j\theta_{1}[t]},e^{j\theta_{2}[t]},\cdots,e^{j\theta_{L}[t]}\right]$ is the passive beamforming vector of the IRS, and $\theta_{l}[t]\in[0,2\pi)$ is the reflecting coefficient of the $l$-th IRS element, wherein $l=\{1,2,\cdots,L\}$.
@@
3.3 Communication Model
Let $s_{n}[t]$ and $\omega_{n}[t]\in\mathbb{C}^{M\times 1}$ denote the symbol transmitted by the BS to the $n$-th user and the corresponding active beamforming vector in time slot $t$, respectively. We consider that the symbol are independent and identically distributed with zero mean and unit variance, expressed as $s_{n}\sim\mathcal{N}(0,1)$. Therefore, the BS transmit signal can be represented by
@@ 3.3 Communication Model
where $\varrho_{n}[t]\sim\mathcal{N}(0,\sigma_{n}^{2})$ represents the additive white Gaussian noise characterized by a variance $\sigma^{2}$ in time slot $t$.
@@
$\gamma_{n}[t]=\frac{\left|\widetilde{H}_{n}^{\text{bu}}[t]\omega_{n}[t]\right|^{2}}{\sum_{i=1,i\neq n}^{N}\left|\widetilde{H}_{n}^{\text{bu}}[t]\omega_{i}[t]\right|^{2}+\sigma^{2}}.$ (11)
3.4 Sensing Model
ISAC can use the transmitted communication signals for sensing operations by sharing the spectrum resources [13], thereby achieving efficient resource utilization. Given that the presence of obstacles impacts the sensing performance of the BS, a low-altitude IRS is deployed to reconfigure the sensing link by adjusting the phase of its reflective elements, enabling the direct target sensing services [25]. Thus, the steering vector from the IRS to the target can be expressed as
$a[t]=\left[1,e^{j\frac{2\pi}{\lambda}\sin({\vartheta^{\text{ST}}}[t])},\cdots,e^{j\frac{2\pi(L-1)}{\lambda}\sin({\vartheta^{\text{ST}}}[t])}\right]^{T},$ (12)
where $\vartheta^{\text{ST}}[t]$ is the angle of interest in time slot $t$.
Then, the received signal of the target is denoted as [11]
$\widetilde{\mathbf{x}}[t]=\mathbf{g}^{\rm H}[t]\left(\sum_{n=1}^{N}\omega_{n}[t]s_{n}[t]\right)+\varrho^{\text{ST}}[t],$ (13)
where $\varrho^{\text{ST}}[t]$ is the noise received at the target. Moreover, $\mathbf{g}^{\rm H}[t]=\alpha_{\rm r}[t]a^{\text{H}}[t]\Phi[t]h^{\text{br}}[t]$, and $\alpha_{\rm r}[t]$ is the large-scale fading coefficient [11]. Thus, the total gain towards the target in time slot $t$ is defined as
@@ 3.4 Sensing Model
$=\mathbf{g}^{\text{H}}[t]\left(\sum_{n=1}^{N}\omega_{n}[t]\omega_{n}^{\text{H}}[t]\right)\mathbf{g}[t].$
Notably, the core task of target sensing is typically to extract the state information of the target from the reflected signal. In the passive target scenario, since the target does not actively transmit signals, its signal reflection ability mainly depends on the received signal power. Therefore, to quantify the sensing ability of the system in a specific direction, we define the sensing rate metric to measure how effectively the system can project useful information toward the target for sensing, which is expressed as [13]
$R^{\text{ST}}[t]=\log_{2}(1+\frac{\mathbf{G}[t]}{\sigma^{2}}).$ (15)
3.5 UAV Mobility and Energy Model
We assume that the UAV flies at a fixed altitude $z^{\text{r}}$ within the target area for task execution. The horizontal location $\left(x^{\text{r}}[t],y^{\text{r}}[t]\right)$ of the UAV (i.e., low-altitude IRS) in time slot $t$ depends on the flight speed $v_{\text{u}}[t]$ and yaw angle $\theta_{\text{u}}[t]\in[-\pi,\pi]$, which can be denoted as
@@ 3.5 UAV Mobility and Energy Model
Due to the limited energy resources of the UAV, optimizing propulsion energy consumption can enhance endurance, thereby extending mission execution time and improving the overall operational efficiency of the system. The propulsion energy of the UAV in time slot $t$ can be expressed as [36, 37]
@@
4 Problem Formulation and Analyses
In this section, we aim to formulate CSEMOP that focuses on the communication, sensing, and energy efficiency performance of the considered low-altitude IRS-assisted ISAC system. We first introduce the optimization variables and objectives. Based on this, the optimization problem is formulated and analyzed.
4.1 Problem Formulation
@@
Objective 2: Sensing rate is a key metric for evaluating the sensing performance, which directly decides the quality and reliability of sensing tasks. Thus, the second objective aims to maximize the sensing rate of the target, which is given by
@@
Objective 3: A UAV is an energy-constrained airborne platform, and reducing its energy consumption can extend operating time of the considered system. Therefore, the third objective focuses on minimizing propulsion energy, expressed as follows:
@@
$-\pi\leq\theta_{\rm u}[t]<\pi,\quad\forall t\in\mathcal{T},$ (21e)
@@
where constraints (21b) and (21c) represent the boundaries of the area within which the UAV performs the mission. Moreover, the flight speed and yaw angle in each time slot are bounded by constraints (21d) and (21e). In addition, constraint (21f) indicates that the transmission power of the BS is less than $P_{\rm max}$. Furthermore, constraint (21g) specifies the phase shift constraint of the IRS.
4.2 Problem Properties
We can find that the formulated CSEMOP has the following characteristics:
Non-convexity: According to the three optimization objectives presented in Eqs. (18)-(20), it can be observed that the optimization variables $\mathbf{\Omega}$, $\mathbf{\Psi}$, $\mathbf{\mathcal{V}}$, and $\mathbf{\Theta}$ are coupled. Moreover, IRS phase shift $\theta_{l}$[t] appears exponentially in the IRS beamforming vector [17]. As a result, CSEMOP is non-convex, making it difficult to solve.
@@ 4.2 Problem Properties
Trade-off: It can be observed that there are conflicts among the three optimization objectives. For example, focusing on the communication performance between the BS and the users will impact the sensing capability of the BS for the target under given channel conditions. Moreover, improving both sensing and communication performance of the system requires rapid deployment of the low-altitude IRS to suitable regions, which increases the UAV propulsion energy consumption. Therefore, there is a trade-off among the three optimization objectives.
5 Proposed Algorithm
In this section, we first clarify the motivation for using DRL to solve CSEMOP. Subsequently, we reformulate the CSEMOP as an MDP. Next, we present the core principles of the conventional DDPG algorithm and the improved factors. Finally, we show the proposed GDMDDPG algorithm and analyze the computation complexity.
5.1 Motivation of Using DRL
For conventional optimization algorithms, the optimization problem is typically decomposed into several subproblems and solved by using the alternating optimization algorithms, which may influence the accuracy of the final result. Moreover, low-altitude IRS makes the system highly dynamic, and the prior knowledge required by traditional optimization algorithms is difficult to obtain in such a dynamic environment. In addition, the considered system evolves over time, making it a significant challenge to balance short-term and long-term gains of the system during task execution. However, traditional optimization algorithms have limitations in achieving such balance.
In contrast, DRL algorithms use the trial-and-error learning mechanism through interactions between the agent and the environment, and possess superior capabilities for balancing short-term and long-term rewards in the dynamic environment, providing a feasible solution to the aforementioned challenges. Therefore, we propose a GDMDDPG algorithm to solve the considered CSEMOP.
5.2 MDP Construction
@@
5.2.1 State Space $\mathcal{S}$
The environment state $s[t]\in\mathcal{S}$ observed by the agent in time slot $t$ is defined as $s[t]=\{t,q^{r}[t],q^{\rm u}[t],q^{\rm g}[t]\}$, where $t$ represents the index of current time slot in the pre-defined time slot set $\mathcal{T}$. Moreover, $q^{r}[t]$ and $q^{\rm g}[t]$ are the locations of the low-altitude IRS and the target in time slot $t$. In addition, $q^{\rm u}[t]=\{q_{1}^{\rm u}[t],q_{2}^{\rm u}[t],\ldots,q_{N}^{\rm u}[t]\}$ is the location set of all users in time slot $t$.
5.2.2 Action Space $\mathcal{A}$
According to the observed environment state $s[t]$ in time slot $t$, the agent can take the action $a[t]\in\mathcal{A}$, which is denoted as $a[t]=\{\omega[t],\phi[t],v_{u}[t],\theta_{u}[t]\}$. Specifically, $\omega[t]=\{\omega_{n}[t]|n\in\mathcal{N}\}$ represents the beamforming matrix of the BS for all users, and $\phi[t]$ represents the phase shifts of the low-altitude IRS in time slot $t$. Moreover, $v_{u}[t]$ and $\theta_{u}[t]$ denote the flight velocity and yaw angle of the UAV in time slot $t$, respectively.
5.2.3 Reward Function $\mathcal{R}$
@@
5.3 Principles of DDPG Algorithm
@@
5.3.1 Actor-Critic Framework
@@
5.3.2 Experience Replay Buffer
@@ 5.3.3 Network Update
To avoid significant changes in the target network during training, which could lead to instability in the agent learning process, the target networks are updated using a soft-update method as follows
@@ 5.4 GDMDDPG
5.4.1 The Limitations of Conventional DDPG
@@
First, in conventional DDPG algorithm, the actor network typically employs MLP as the core architecture due to the simplicity and ease of implementation. However, in more complex environments, the fixed MLP architecture may fails to fully capture and analyze the intricate information present in the environment states, which in turn limits the decision-making capabilities of the agent [40]. Notably, diffusion model provides a viable solution to address the above limitations due to its strong ability in modeling complex data distributions and generating high-quality samples [41]. By replacing the traditional MLP with the diffusion model, the actor network can more effectively analyze the environment states, thereby generating superior actions and further enhancing the overall system performance.
@@
Second, DDPG is based on deterministic policy, where the actor network directly outputs a specific action value rather than a probability distribution over actions. Such a deterministic policy limits the agent ability to explore the action space effectively [42]. To address this issue, noise can be added to the actions generated by the actor network, introducing randomness to promote the agent exploration of the action space. Moreover, this noise perturbation mechanism helps alleviate the tendency of the DDPG to output action boundary values during early training phase, thereby enhancing the convergence speed of the training process.
@@
Finally, the experience replay buffer in conventional DDPG employs a random sampling strategy, which might fail to prioritize experiences that hold significant learning value, resulting in slower convergence [43]. To address this issue, the RPER sampling mechanism, which integrates ERE and PER techniques, can be incorporated into the network update process of the DDPG. Specifically, ERE technique ensures that more recent experiences are sampled more actively. Based on this, the PER technique assigns a priority to each experience sample, allowing those that contribute more to learning to be sampled with higher probability. As such, combining PER and ERE can effectively mitigate the limitations of the random sampling strategy.
@@
In the following, we present detailed explanation of the diffusion model-based actor network, noise perturbation mechanism, and RPER sampling mechanism.
5.4.2 Diffusion Model-based Actor Network